References Review View Help Open in Desktop App Tell me what you want to do Editing 11 v Dictate Α Α Β Ι Ο ν Αν Α. E 日 A e
Answer #1

Computation of number of shares, stock dividend, value of stock dividend

            Number of shares                   = 500000+400000 = 900000

            Stock dividend                        = 10%*900000 = 90000 shares

            Value of stock dividend         = 90000*8 = 72000

Computation of big auto part’s total shareholder’s equity December 31, 2019

Particulars

amount

Common stock capital

500000 shares at $7 each

3500000

400000 share at $8.5 each

3400000

Add: stock dividend 90000 shares at $8 each

720000

A =

7620000

big auto part inc retained earnings

Net income

920000

Less: cash dividend

140000

780000

Less: stock dividend (90000 shares@ $8 each)

720000

B =

60000

            Big auto part’s total shareholder’s equity December 31, 2019 = (A+B)

                        = 7620000+60000 = 7680000